# Coherent Load Profile Synthesis with Conditional Diffusion for LV Distribution Network Scenario Generation

### v1

@@ 4 Synthesis Results
Table 1: Error metrics for each model. Results are calculated on the entire unseen test set using scaled values. Best scores are highlighted in bold.
Table 1 calculates the score of each metric for each model. When comparing unconditional models, LVGenU outperforms the GMM and WGAN in synthesis and reconstruction with lower MSE and MMD scores. GMM and WGAN do exhibit a lower Wasserstein distance, with WGAN further demonstrating a significantly lower marginal score. The LVGenU is trained on MSE and consequently does not pay direct attention to the distribution shape. On the contrary, the WGAN contains the Wasserstein distance in its loss function and hence the model is directly optimised to capture the distribution shape without considering the temporal coherence of the samples (as it will be discussed later when examining Autocorrelation Function performance; Figure 7). Generative models that utilise conditions in LVGenWC and LVGenWCS significantly improve all these scores, with an extremely good match between the synthesised and real profiles.
@@
Figure 5 plots the active and reactive power distribution curves of the real dataset and each of the proposed models. Results show a similar trend to what is observed in the Marginal score and Wasserstein Distance; the increase in conditions results in a more accurate capture of the distribution. The GMM provides a good capture of the overall distribution for the active and reactive power; however, it fails to capture both tails of the distribution. This is due to the GMM being unable to capture the higher-order moments. The WGAN demonstrates a better overall fit to the distributions compared to the GMM, mainly due to its improved capture of the peak of each distribution. However, it still fails to accurately capture the lower tail. This behaviour reflects the influence of the WGAN’s penalty term, which contributes to its improved Wasserstein Distance and Marginal Score. However, this emphasis results in weaker performance in all other metrics, namely MSE and MMD. LVGenU demonstrates a pessimistic generation rarely diverging from the median samples in order to minimise the MSE loss function. The lower MSE allows the model to achieve better reconstruction of load profiles; however, this comes at the cost of a less accurate distribution capture, highlighted by a higher Marginal score and Wasserstein distance compared to the GMM and WGAN. LVGenWC improves upon LVGenU by generating a wider variety of samples and achieving a more accurate capture of the distribution, as demonstrated by an improvement in each of the distance-based metrics. This is due to the distribution containing a less significant oversample of the median, and not as severe a miss of the extreme values. However, both issues persist, though not to the same extent as with LVGenU. The distribution generated by LVGenWCS (orange) almost perfectly matches the real distribution (red), making it difficult to distinguish between the two in the figures, as LVGenWCS closely overlaps the real distribution. The model is also able to capture multi-modal behaviour in the tail of the distribution for the reactive power. This leads to a substantially lower MMD score compared to all other models, representing the most significant improvement in the metric by a wide margin.
@@
Figure 6 contains the active and reactive power decile plots for the real dataset and each of the proposed models. The figure plots the typical load profiles throughout various deciles for active and reactive power. Results show the GMM misses the real load profile at every decile for active and reactive power, all to a similar degree; these consistent misses result in a high MSE. WGAN is in the correct range for the 90th decile, but misses the shape of the load profile. In the other deciles, WGAN misses more severely than the GMM, resulting in an even higher MSE due to these extreme missed values. LVGenU is much closer to the real data for active and reactive power at the 10th decile compared to GMM and WGAN. At the 50th decile, similar results to the GMM are obtained. The model completely misses the 90th decile, however still manages to retain a lower MSE when compared to the benchmark models. The generally high MSE for each unconditional model results from their uninformed generation process, as neither model has any indication of the load profile it is attempting to reconstruct (guidance that subsequent models receive through their conditions). LVGenWC provides a more accurate capture of active and reactive power at the 10th and 50th deciles, performing comparably to LVGenWCS in these areas. However, it struggles to capture the 90th decile, a trend similar to what was observed in Figure 5. This is reflected in the MSE, where LVGenWC shows improvement due to its better capture of the 10th and 50th deciles, along with a less severe miss at the 90th decile. LVGenWCS further improves the MSE by achieving a much better capture of the 90th decile for both active and reactive power.
@@
Together, the metrics and visualisations outline the performance of each model and demonstrate the improvement from the baseline GMM and WGAN to a diffusion model in an unconditional perspective. This is through the ability of the diffusion model to capture higher-order moments in the data and better replicate the temporal structure of real data. Then the addition of conditional inputs led to further improvements being observed in the results. The LVGenWC model shows promise as a scenario generation tool, but does not adequately sample from the extremes of the distribution, nor does it consistently recreate load profiles to high accuracy. To address this limitation, additional inputs or modifications to the model may be required to capture these extreme values effectively. However, for the specific goal of this study, LVGenWCS is far more suitable as it can generate specific load profiles from individual substations to a high quality and accuracy right across the distribution. In subsequent sections, it is demonstrated how the models generated aggregated load profiles perform in a wider power systems context through the application of power systems models. Although informative in terms of model performance, synthesis metrics alone are not sufficient. A model must also produce robust phase angles and voltage estimates across the entire network to serve as a credible replacement for rigorous monitoring.
@@ 5.2 Load Flow Setup and Analysis
The load flow analysis is used to ensure that the power system operates within its network limits of 0.05 p.u. for voltage magnitude and 10 degrees for phase angle, which are standard values in GB distribution networks. Thus, determining the most economical generation dispatch [43]. The load flow used is based on the Newton-Raphson method and is performed using the Python library pandapower [44]. Given an $n$-bus system with one slack bus, which has a constant 1.0 p.u. voltage magnitude and zero phase angle. The complex power at the node $n$ can be calculated as [45]:
@@ 5.3 Load Flow Simulation Results
Figure 9: Comparison scatterplot of Voltage Magnitude predictions versus ground truth for each model.
@@
Figure 9 and Table 2 present a comparison of the voltage magnitude results obtained from the benchmark and proposed diffusion models. Models tend to perform well in estimating higher voltage magnitudes when the system load is at a minimum, but tend to be less accurate at lower voltage magnitudes when the load is near its peak. In general, most methods overestimate the voltage magnitude, which could result in inappropriate decisions by system operators. In contrast, the LVGenWCS method consistently captures the trend of the voltage magnitude accurately, regardless of whether the values are high or low.
@@
Figure 10: Single LV bus time series of hourly average voltage magnitude demonstrating temporal fidelity between real and LVGenWCS data with min and max thresholds.
Furthermore, Figure 10 compares the LVGenWCS with the real data for a single bus test case. The LVGenWCS model was run ten times, and the results were averaged to produce a representative output, with confidence intervals also reported based on the model’s upper and lower outputs at each timestep. The findings demonstrate that the LVGenWCS model provides reliable results that replicate the dynamic behaviour of the network. The LVGenWCS model closely aligns with the real load data, accurately capturing voltage patterns with reduced deviations in amplitude. Although some of the troughs are overestimated in certain model results, the average of the results is sufficiently accurate, with an acceptable error margin. This indicates that the LVGenWCS model offers reliable accuracy for the network and serves as a good fit for predicting voltage magnitudes.
Figure 11: Comparison scatterplot of Phase Angle predictions versus ground truth for each model.
@@
The results in Table 3 also show that the LVGenWCS method outperforms the other methods, again demonstrating MAE that is roughly five times lower than the next best model, and ten times lower for the 5th percentile. The reactive power values here are considerably low due to the data only consisting of resident loads. The results indicate that only the LVGenWCS method shows potential for wider application in reactive power prediction for networks containing commercial or industrial loads, thereby providing system operators with more reliable day-ahead predictions.
@@
Figure 12: Single LV bus time series of hourly average phase angle demonstrating temporal fidelity between real and LVGenWCS data with min and max thresholds.
@@
Overall, the strong correspondence between the real data and LVGenWCS model outputs for both voltage magnitude and phase angle highlights the effectiveness of the LVGenWCS model in replicating complex voltage characteristics based on the load profile. This high level of accuracy suggests that the model can be reliably used for both prediction and analysis in distribution power systems.
6 Conclusion
Transitioning towards Distribution System Operation in legacy power systems requires understanding of load interactions at the sub-regional level to avoid thermal and voltage constraint violations. Extensive monitoring would ultimately inform mitigating strategies to a high standard, but it is prohibitively expensive to deploy and maintain comprehensively across distribution networks owing to the sheer volume of assets. While load behaviour synthesis offers a convenient alternative, demand characteristics at low voltage substations must be realistically diverse so that higher voltage level behaviour is accurately represented: lack of diversity results in voltage collapses and thermal exceedances, excessive diversity will excessively smooth out any extremes. To address the problem of realistic distribution load profile synthesis, this work has contributed a Generative Conditional Diffusion Model for reconstructing LV substation load profiles based on a minimum number of cues. Realism was demonstrated on an individual level through the validation of temporal and statistical characteristics, while coherence in the wider power system context was demonstrated through the propagation of synthesised LV loads through a representative MV network. The load flow results demonstrate that the synthesised LV substation profiles have a similar impact on the network when compared with results obtained from using meter data. The implication of the equivalent load flow result is that diversity across LV substations synthesised by the diffusion model has temporal variability, which reflects load diversity appropriately in most cases. This model provides a good synthesis of the substation base load profiles, which is necessary for power system studies and is a prerequisite to benchmarking the impacts of LCT penetration. Future development will entail retraining the model with reliable LCT metadata, such as the proportions of the distributed energy sources adopted under the LV substation, such as photovoltaic (PV), wind and EV charging, to better learn and replicate the effects of these diversity changing loads. From a behavioural perspective, understanding rebound effects from time of use tariffs (for instance, extrapolating [20]), there is also the potential for model trust to be questioned, which motivates the need for future work in model explainability to account for the plausibility of what has been captured by the model and how it produces subsequent load representations on which planning and operational decisions are made.

### v2

@@ 4 Synthesis Results
Tail-Focused MSE 7.61 7.56 7.70 4.32 0.65
Table 1: Error metrics for each model. Results are calculated on the entire unseen test set using scaled values. The results represent the average of each metric’s performance calculated separately on active and reactive power, respectively. Some metrics contain the $\pm$ standard deviation from the mean on a per-substation basis. Best scores are highlighted in bold.
Table 1 reports the values of each evaluation metric for all models considered. The unconditional models are optimised according to fundamentally different objective functions, which directly explains their relative performance across metrics. LVGenU is trained to minimise MSE at the level of individual substation load profiles, prioritising accurate reconstruction and temporal coherence. In contrast, the GMM is optimised via expectation–maximisation to maximise the log-likelihood of the aggregated load distribution across all substations, while the WGAN is directly optimised to minimise the Wasserstein distance between aggregated load distributions. As a consequence, each model performs most strongly on metrics that align with its underlying optimisation objective.
Consistent with this, LVGenU achieves lower values for MSE, MMD, and MiVo relative to the GMM and WGAN baselines. The reduced MSE indicates improved reconstruction fidelity at the substation level, while the lower MMD reflects closer alignment between the joint distribution of generated active and reactive power profiles and the corresponding real data. The lower MiVo score further indicates improved preservation of variability and co-dependence between active and reactive power across individual substations. Together, these metrics demonstrate that LVGenU more effectively preserves the temporal structure of individual load profiles, which is essential for maintaining realistic voltage trajectories and dependencies between active and reactive power, a trend that is also reflected in the autocorrelation function shown in Figure 7.
At the same time, LVGenU does not outperform the other unconditional baselines on metrics that primarily assess distributional shape, including the Marginal Score, Wasserstein distance, and Tail-Focused MSE. In these cases, the GMM and WGAN can achieve lower scores, reflecting their explicit optimisation for aggregated distributional similarity. In particular, the WGAN incorporates the Wasserstein distance directly into its loss function, enabling improved global distribution capture at the expense of substation-level temporal realism and coherence. This highlights a clear trade-off between capturing global distributional shape and preserving substation-level temporal realism.
Overall, LVGenU does not capture the distribution of the aggregated load profiles but exhibits strong temporal coherence and reconstruction fidelity. These characteristics are particularly important for realistic LV load synthesis, where preserving substation-level dynamics is more critical than matching aggregated distributional shape. However, misrepresentation of extreme operating conditions may lead to an underestimation of peak loads, which in a power systems context could result in the failure to model critical scenarios such as thermal excursions. These limitations motivate the extension of LVGenU through additional conditioning in the proposed framework.”
The introduction of conditions in the further models results in a significant improvement in all metrics and, in turn, mitigates the issues in the LVGenU model. LVGenWC is able to build on the performance of LVGenU, improving in all metrics. The conditions not only improved on the struggling areas of LVGenU, such as Wasserstein distance and Tail Capture, but also significantly improved on areas where it was already outperforming the other baseline models. The result of these improvements is plausible substation loads with improved representation of the distribution extremes. LVGenWCS then improves all metrics further to a level where load reconstruction can be carried out with high accuracy.
@@
Figure 5 plots the active and reactive power distribution curves of the real dataset and each of the proposed models. Results show a similar trend to what is observed in the Marginal score and Wasserstein Distance; the increase in conditions results in a more accurate capture of the distribution. The GMM provides a good capture of the overall distribution for the active and reactive power; however, it fails to capture both tails of the distribution. This is due to the GMM’s fitted mixture components not adequately capturing low‑probability regions in the unseen test set. The WGAN demonstrates a better overall fit to the distributions compared to the GMM, mainly due to its improved capture of the peak of each distribution. However, it still fails to accurately capture the lower tail. This behaviour reflects the influence of the WGAN’s penalty term, which contributes to its improved Wasserstein Distance and Marginal Score. However, this emphasis results in weaker performance in all other metrics, namely MSE and MMD. LVGenU demonstrates a pessimistic generation rarely diverging from the median samples in order to minimise the MSE loss function. The lower MSE allows the model to achieve better reconstruction of load profiles; however, this comes at the cost of a less accurate distribution capture, highlighted by a higher Marginal score and Wasserstein distance compared to the GMM and WGAN. However, despite this limitation, the lower MMD score achieved by LVGenU compared to the other unconditional baselines indicates that, on a per-substation basis, LVGenU aligns more closely with the real data. The tail-focused MSE metrics shows although models exhibit differences of distributional capture at extremes, their point-wise errors are largely similar. LVGenWC improves upon LVGenU by generating a wider variety of samples and achieving a more accurate capture of the distribution, as demonstrated by an improvement in each of the distance-based metrics and reduced Tail-Focused MSE. This is due to the distribution containing a less significant oversample of the median, and not as severe a miss of the extreme values. However, both issues persist, though not to the same extent as with LVGenU. The distribution generated by LVGenWCS (orange) almost perfectly matches the real distribution (red), specifically at the upper distribution tail, which previous models failed to capture. This is shown by a drastic reduction in Tail-Focused MSE. These results make it difficult to distinguish between the true and generated distributions in the figures, as LVGenWCS closely overlaps the real. The model is also able to capture multi-modal behaviour in the tail of the distribution for the reactive power. This leads to a substantially lower MMD score compared to all other models, representing the most significant improvement in the metric by a wide margin.
@@
Figure 6 contains the active and reactive power decile plots for the real dataset and each of the proposed models. The figure plots the typical load profiles throughout various deciles for active and reactive power. Results show the GMM misses the real load profile at every decile for active and reactive power, all to a similar degree; these consistent misses result in a high MSE. WGAN is in the correct range for the 90th decile, but misses the shape of the load profile. In the other deciles, WGAN misses more severely than the GMM, resulting in an even higher MSE due to these extreme missed values. LVGenU is much closer to the real data for active and reactive power at the 10th decile compared to GMM and WGAN. At the 50th decile, similar results to the GMM are obtained. The model completely misses the 90th decile, however still manages to retain a lower MSE when compared to the benchmark models. The generally high MSE for each unconditional model results from their uninformed generation process, as neither model has any indication of the load profile it is attempting to reconstruct (guidance that subsequent models receive through their conditions). Each model’s limited ability to capture behaviour in the upper quantiles is directly correlated with the Tail-Focused MSE. LVGenWC provides a more accurate capture of active and reactive power at the 10th and 50th deciles, performing comparably to LVGenWCS in these areas. However, it struggles to capture the 90th decile, a trend similar to what was observed in Figure 5. This is reflected in both MSE metrics, where LVGenWC shows improvement due to its better capture of the 10th and 50th deciles, along with a less severe miss at the 90th decile. LVGenWCS further improves the MSE metrics by achieving a much better capture of the 90th decile for both active and reactive power.
@@
Together, the metrics and visualisations outline the performance of each model and demonstrate the improvement from the baseline GMM and WGAN to a diffusion model in an unconditional perspective. This is through the ability of the diffusion model to capture higher-order moments in the data and better replicate the temporal structure of real data. Then the addition of conditional inputs led to further improvements being observed in the results. The LVGenWC model shows promise, being capable of generating substations from a much greater range of the distribution, accounting for many of the substation loads in the dataset. However, it does not adequately sample from the extremes of the distribution, nor does it consistently recreate load profiles to high accuracy. LVGenWCS is far more suitable as it can generate specific load profiles from individual substations with high quality and accuracy right across the distribution. The results show that the daily statistics were especially required in order to accurately model the extremes of the distribution; substation metadata alone is not enough. While monitoring data is limited targetting specific substations based on geography or other possible factors may provide a potential solution. In subsequent sections, it is demonstrated how the models generated aggregated load profiles perform in a wider power systems context through the application of power systems models. Although informative in terms of model performance, synthesis metrics alone are not sufficient. A model must also produce robust phase angles and voltage estimates across the entire network to serve as a credible replacement for rigorous monitoring.
@@ 5.2 Load Flow Setup and Analysis
The load flow analysis is used to verify that the power system operates within network limits ($\pm 0.05\,\text{p.u.}$ for voltage magnitude and 10 degrees for phase angle), which are standard values in GB distribution networks. This enables the identification of constraint violations and assessment of system feasibility. The load flow used is based on the Newton-Raphson method and is performed using the Python library pandapower [39]. Given an $n$-bus system with one slack bus, which has a constant 1.0 p.u. voltage magnitude and zero phase angle. The complex power at the node $n$ can be calculated as [40]:
@@ 5.3 Load Flow Simulation Results
Figure 9: Comparison scatterplot of Voltage Magnitude predictions versus ground truth for each model. Each panel contains acceptance bands representing acceptable deviations from the true values ($\pm 1\%$ for voltage magnitude). Points within the acceptable band are plotted in black, while points outside are plotted in orange.
@@
Figure 9 and Table 2 present a comparison of the voltage magnitude results obtained from the benchmark and proposed diffusion models. To ensure statistical robustness and account for the stochastic nature of the generative process, all LVGen models were evaluated over 10 independent runs; the results reported herein represent the average values across these runs. In Figure 9, the red reference line represents perfect agreement between the true and predicted values. The black dashed lines define an acceptance band corresponding to $\pm 1\%$ of the voltage magnitude, in accordance with statutory limits. Points within the band are plotted in black, while points outside the band are highlighted in orange.
The models generally perform well in estimating higher voltage magnitudes when the system load is at a minimum, but tend to be less accurate at lower voltage magnitudes when the load is near its peak. In general, most methods overestimate the voltage magnitude, which could result in inappropriate decisions by system operators. In contrast, the LVGenWCS method consistently captures the trend of the voltage magnitude accurately, regardless of whether the values are high or low.
@@
Figure 10: Single LV bus time series of hourly average voltage magnitude demonstrating temporal fidelity between real and LVGenWCS data with min and max thresholds. Variability in model output over multiple stochastic generations is highlighted using min-max error bars.
Furthermore, Figure 10 compares the LVGenWCS with the real data for a single bus test case. The LVGenWCS model results are shown with confidence intervals also reported based on the model’s upper and lower outputs at each timestep. The findings demonstrate that the LVGenWCS model provides reliable results that replicate the dynamic behaviour of the network. The LVGenWCS model closely aligns with the real load data, accurately capturing voltage patterns with reduced deviations in amplitude. Although some of the troughs are overestimated in certain model results, the average of the results is sufficiently accurate, with an acceptable error margin. This indicates that the LVGenWCS model offers reliable accuracy for the network and serves as a good fit for predicting voltage magnitudes.
Figure 11: Comparison scatterplot of Phase Angle predictions versus ground truth for each model. Each panel contains acceptance bands representing acceptable deviations from the true values ($\pm 5\%$ for phase angle). Points within the acceptable band are plotted in black, while points outside are plotted in orange.
@@
The results in Table 3 also show that the LVGenWCS method outperforms the other methods, again demonstrating an MAE that is roughly five times lower than the next best model, and ten times lower for the 5th percentile. The reactive power values here are considerably low due to the data only consisting of resident loads. The results indicate that only the LVGenWCS method shows potential for wider application in reactive power prediction for networks containing commercial or industrial loads, thereby providing system operators with more reliable day-ahead predictions.
@@
Figure 12: Single LV bus time series of hourly average phase angle demonstrating temporal fidelity between real and LVGenWCS data with min and max thresholds. Variability in model output over multiple stochastic generations is highlighted using min-max error bars.
@@
Overall, improvements in model performance with the introduction of initial conditioning in the LVGenWC model do not extend to the power flow case study. This shortcoming does not translate to the network being generated being unrealistic or not plausible, but rather that it does not resemble the true network in this example. In order to accurately model the network in the load flow analysis, the daily load statistics were required. The strong correspondence between the real data and LVGenWCS model outputs for both voltage magnitude and phase angle highlights the effectiveness of the LVGenWCS model in replicating complex voltage characteristics based on the load profile. This high level of accuracy suggests that the model can be reliably used for both prediction and analysis in distribution power systems. It is expected that access to a wider range of reliable metadata in the future, and its subsequent incorperation to the models can limit the requirement for monitoring to perform accurate power flow analysis.
6 Conclusion and Future Work
Transitioning towards Distribution System Operation in legacy power systems requires an understanding of load interactions at the sub-regional level to avoid thermal and voltage constraint violations. Extensive monitoring would ultimately inform mitigating strategies to a high standard, but it is prohibitively expensive to deploy and maintain comprehensively across distribution networks owing to the sheer volume of assets. While load behaviour synthesis offers a convenient alternative, demand characteristics at low voltage substations must be realistically diverse so that higher voltage level behaviour is accurately represented: lack of diversity results in voltage collapses and thermal exceedances, excessive diversity will excessively smooth out any extremes.
To address the problem of realistic distribution load profile synthesis, this work has contributed various Generative Diffusion Models capable of reconstructing LV substation load profiles based on a range of conditional cues. Realism was demonstrated on an individual level through the validation of temporal and statistical characteristics, while coherence in the wider power system context was demonstrated through the propagation of synthesised LV loads through a representative MV network.
The key outcomes of this study are a fleet of models that contrast unconditional synthesis with progressively conditioned generation to support a range of practical use cases. Results showed that while an unconditional model was capable of generating realistic load profiles, it was not able to synthesise specific load behaviours, making it of limited use from both a scenario generation and reconstruction perspective. The introduction of conditional metadata in LVGenWC, such as calendar information or weather variables that are often readily available, greatly improved synthesis quality and enabled the model to realistically synthesise a far greater range of behaviours representing a significant number of substations. These models provide a synthesis of substation base load profiles that is suitable for downstream power system studies and serve as a prerequisite for benchmarking the impacts of LCT penetration or changes in customer numbers. However, LVGenWC struggled at distribution extremes, causing it to perform poorly in the load flow analysis. While LVGenWC can generate a network of plausible loads, it struggles to recreate real networks due to mismodelling of individual substations, indicating that conditioning on lightweight non-substation specific metadata alone is insufficient to fully capture behaviour at the distribution extremes. The introduction of daily statistics (daily min, mean, and max of load profiles) improved the model’s capture at the distribution extremes, and load flow analysis demonstrated a similar impact on the network to that observed using real meter data. Although such data may not always be available, these models are capable of generating highly accurate networks which could be used for various other tasks within the power system, such as stress testing.
There are a number of areas of limitation which future work aims to address. The results show that although a conditional diffusion model driven by substation metadata and weather is capable of generating plausible univariate substation load profiles, the associated performance gains do not consistently translate into realistic network-level behaviour in the power flow case study. In particular, when substations are generated independently and without synchronisation, additional statistical conditioning is required to reproduce physically meaningful aggregate behaviour. This observation motivates future work aimed at improving the capture of inter-substation co-behaviour and enabling the modelling of LV networks without reliance on extensive monitoring data.
The present study focuses on modelling LV base load and does not explicitly capture the distinct dynamics associated with high-penetration LCTs. This choice reflects the current lack of reliable, large-scale LCT datasets suitable for conditioning deep generative models. As such datasets become available, the proposed framework can be retrained with additional conditional priors to explicitly represent LCT-driven demand dynamics and support deep-decarbonisation scenario analysis. Furthermore, the models are trained and tested on a single DNO dataset, and their evaluation must be extended to datasets from other DNOs and geographical contexts to ensure generalisability in other regions, network topologies and customer mixes.
Moreover, the evaluation primarily assesses model performance under historical operating conditions. Identifying and characterising structural breaks or regime changes in LV load behaviour is a research challenge in its own right and would require a dedicated methodological contribution. While testing on unseen substations provides a partial mitigation of domain shift, explicitly evaluating model robustness under structural change remains an open challenge and a key direction for future work.
Finally, the proposed diffusion-based framework is inherently data-driven and does not explicitly embed physical or first-principles load models. While power-flow-based evaluation demonstrates that the independently synthesised profiles produce physically plausible network behaviour, the internal representations learned by the model remain difficult to interpret. Improving model transparency and explainability, particularly in relation to physically meaningful power-system quantities, is therefore an important direction for future research.